Super Mario is standing on a 2.9 m tall mushroom. When he jumps off of it,what will his velocity be right as he is about to hit the ground ? (Number only rounded to the nearest tenth ex:4.2)


1
Expert's answer
2021-01-13T11:35:52-0500

We can find the velocity of Super Mario from the law of conservation of energy:


"PE=KE,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot 9.8\ \dfrac{m}{s^2}\cdot 2.9\ m}=7.5\ \dfrac{m}{s}."

Answer:

"v=7.5\ \dfrac{m}{s}."